find the area of a rectangle given that its perimeter is 60 millimeters and its width is 12 millimeters
Answer by xbscpex(1)

Perimeter is 60mm.
W= 12mm
2L + 2W = 60 divide both sides by 2
L + W = 30 ==> Equation 1
substitute W in equation 1
L + 12 = 30
L = 18
Solve area
A = L * W
A = 18 * 12
A = 216 mm^2